OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (imtiaz7):

it's (b) y+1=4(x+3)

OpenStudy (imtiaz7):

By using the formula: Using the point slope form: y-y1=m(x-x1) and here x1 and y1 are -3 and -1 and slope is 4

OpenStudy (anonymous):

thank you!